OpenStudy (anonymous):

ABC is a triangular flag with BC = (5x-2) cm. DEF is a triangular flag with FE = (x+6) cm. The measure of angle BAC is equal to the measure of angle FDE. The measure of angle ABC is equal to the measure of angle DEF. What is the length of the side BC of triangle ABC? 2 cm 3 cm 5 cm 8 cm

OpenStudy (anonymous):

okay @Katt♥ we start by drawing a simple figure to explain what's going on  now you can use the sin law to get this:\[\frac{ \sin \theta1 }{ x+6 }=\frac{ \sin \theta1 }{ 5x-2}\] immediately you realize that 5x-2 = x+6 so x=2